solve x by factoring
3+5x-2x^2=0

The answer is2, -1/2
I worked this one but not sure if my working is correct could you please work it out for me....thanks!

Answer by scott8148(6628)

you need to combine the factors of 3 and -2 to get 5 ... like maybe 2*3 and 1*-1

(3-x)(1+2x)=0 ... x=-1/2 and x=3 (not 2)